OpenStudy (anonymous):

Write a quadratic equation having the given numbers as solutions. -9 and -7

OpenStudy (saifoo.khan):

x = -9 (x + 9) x = -7 (x + 7) Multiply using FOIL (x+7)(x + 9)

OpenStudy (anonymous):

that is the answer? or is it x^2+16=0??

OpenStudy (saifoo.khan):

x^2+16 x+63

OpenStudy (anonymous):

I got that thanks

OpenStudy (saifoo.khan):

welks.

OpenStudy (anonymous):

there is one more I need help with

OpenStudy (saifoo.khan):

Sure, post that on left. :)